A shot-putter throws the shot with an initial speed of 11.2 m/s from a height of 5.00 ft above the ground. What is the range of the shot if the launch angle is (a) 18.0 ∘ , (b) 35.0 ∘ , (c) 43.0 ∘ ?
Answer #1

We know

vertical component of velocity=vsin(theta)=11.2sin(theta)

height=h=5 ft=1.524 m

putting

h=1.524=-11.2sin(theta)*t+4.9t^2

for

(a) theta=18,

h=1.524=-11.2sin(18)*t+4.9t^2

t=1.013

range=11.2*cos(18)*t=10.79 m

(b)

theta=35,

h=1.524=-11.2sin(35)*t+4.9t^2

t=1.515

range=11.2*cos(35)*t=13.9 m

(c)

theta=43,

h=1.524=-11.2sin(43)*t+4.9t^2

t=1.738

range=11.2*cos(43)*t=14.24 m